Please help me answer this question:
Does 3b-2a equal 3b-2a/1? Please Explain your reasoning.
Thank you

Answer by TimothyLamb(4379)

x = x/1
x*1 = x*1/1
x*1 = x
x = x
---
if the question is:
does (3b-2a) equal (3b-2a)/1 ...
---
let x = (3b-2a):
and given that:
x = x/1
we can say that:
(3b-2a) = (3b-2a)/1
---
also, if the question is:
does 3b-2a equal 3b-(2a)/1 ...
---
let x = (2a):
and given that:
x = x/1
we can say that:
(2a) = (2a)/1
and we can say that:
3b-2a = 3b-(2a)/1
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---